# Reinforcement Learning with Verifiable Rewards: GRPO’s Effective Loss, Dynamics, and Success Amplification

### v3

@@ Abstract.
Group Relative Policy Optimization (GRPO) was introduced in (Shao et al., 2024) and used successfully to train DeepSeek-R1 (Guo et al., 2025) models for promoting reasoning capabilities of LLMs using verifiable or binary rewards. We show in this paper that GRPO with verifiable rewards can be written as a Kullback–Leibler ($\mathsf{KL}$) regularized contrastive loss, where the contrastive samples are synthetic data sampled from the old policy. The optimal GRPO policy $\pi_{n}$ can be expressed explicitly in terms of the binary reward, as well as the first- and second-order statistics of the old policy ($\pi_{n-1}$) and the reference policy $\pi_{\mathrm{ref}}$. Iterating this scheme, we obtain a sequence of policies $\pi_{n}$ for which we can quantify the probability of success $p_{n}$. We show that the probability of success of the policy satisfies a recurrence that converges to a fixed point of a function that depends on the initial probability of success $p_{\mathrm{ref}}$ and the regularization parameter $\beta$ of the $\mathsf{KL}$ regularizer. We show that the fixed point $p^{*}$ is guaranteed to be larger than $p_{\mathrm{ref}}$, thereby demonstrating that GRPO effectively amplifies the probability of success of the policy.
@@ 1. Introduction
We focus in this paper on Reinforcement Learning with Verifiable Rewards (RLVR) using GRPO, as recently termed by (Lambert et al., 2024). Following (Lambert et al., 2024), we distinguish three types of verifiable rewards in the context of LLM training:
  1. (1)
    Correctness Verification. This corresponds to a binary reward that can be obtained via string matching between the generated response and a gold-standard answer—if such an answer exists—for example, in math problems with known solutions. This type of reward has been used in (Guo et al., 2025) and subsequently in open-source implementations such as Open-R1 (Hugging Face, 2024) and DeepScaleR (Luo et al., 2025). When a gold-standard answer does not exist, one can resort to an LLM as a judge to assess the correctness of the response within the training loop, as done in deliberative alignment (Guan et al., 2025).
  2. (2)
    Verification via Execution. In code generation, a code interpreter is used to execute the generated code, producing a 0/1 reward for fail/pass. A battery of unit tests can also be executed to verify the correctness of the code, resulting in a binary reward. Open-R1 (Hugging Face, 2024) recently open-sourced this type of reward evaluation.
  3. (3)
    Verifiable Constraints. Finally, formatting constraints on the output or refusals to answer can be enforced using simple binary rewards to guide RL training for LLMs (Guo et al., 2025) (Lambert et al., 2024).
Verifiable reward balance simplicity and bias and are thought to be less prone to reward hacking than reward models learned from preference data . Reward hacking is a common issue in reinforcement learning where the policy learns to over-optimize a reward leading to a lower quality of the model (Gao et al., 2023). While verifiable rewards are more resilient to reward hacking, Lambert et al. (2024) showed that for low regularization of the $\mathsf{KL}$ constraint to the reference model, reward hacking occurs when using verifiable constraints. Hence we study in this paper $\mathsf{KL}$-regularized Reinforcement Learning with Verifiable Rewards using GRPO. We note that a recent paper (Vojnovic and Yun, 2025) studies GRPO with a focus on the policy obtained using an approximation of the $\mathsf{KL}$ divergence used in practical implementations.
@@
    We show in Section 2 that GRPO with verifiable Rewards can be cast as an adaptive weighted Contrastive Loss between samples from the old policy with 0/1 rewards.
@@
    Armed with this contrastive loss interpretation, we show in Section 3 that GRPO dynamics as the old policy is replaced with the current optimal policy, result in a closed form recursion for the optimal policy, where the optimal policy $\pi_{n}$ can be expressed in terms of the reference policy $\pi_{\mathrm{ref}}$, the old policy $\pi_{n-1}$ and the probability of success of the old policy $p_{n-1}$ (the frequency of reward “1” of generated responses for a given prompt).
@@
    Computing the probability of success under $\pi_{n}$, we show in Section 3 that it satisfies a recursion in time, leading to a fixed point equation. We show in Section 4 that under mild assumptions GRPO’s probability success converges asymptotically to this fixed point solution.
  4. (4)
    We show in Section 4 that this fixed point probability of success is guaranteed to be larger than the probability of success of the reference model, proving thereby that GRPO indeed amplifies the probability of success as observed experimentally.
  5. (5)
    Finally we show in Section 5 that for approximate policies obtained for instance by gradient descent, the probability of success remains close to the fixed point probability of success as long as the approximation, statistical and optimization errors remain small.
@@ 2. GRPO With verifiable Rewards as an Adaptive Weighted Contrastive Loss
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)-\beta\mathsf{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$ (GRPO-Clip)
where the advantage for an outcome $o$, $A(q,o)$ is given by:
$A(q,o)=\frac{r(q,o)-\mathbb{E}_{o^{\prime}\sim\pi_{\theta_{\text{old}}}(.|q)}r(q,o^{\prime})}{\sqrt{\mathsf{Var}_{o^{\prime}\sim\pi_{\theta_{\text{old}}}(.|q)}r(q,o^{\prime}))}},$ (1)
where $\mathrm{Var}$ is the variance and for $\epsilon\in[0,1]$: $f_{\epsilon}(x,y)=\min(xy,\text{clip}(x,1-\epsilon,1+\epsilon)y).$
To simplify Equation GRPO-Clip, let us consider this objective without the clipping ($\epsilon\to+\infty$) we obtain:
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}A(q,o)-\beta\mathsf{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$ (GRPO)
@@
We see that GRPO optimizes the whitened reward (the advantage $A(q,o)$) using importance sampling from the “old” policy while maintaining the optimized policy close to $\pi_{\mathrm{ref}}$ as measured by the $\mathsf{KL}$ divergence. If furthermore the clipping is used as in (GRPO-Clip), the likelihood ratio between the policy and the old policy is maintained within a range $[1-\epsilon,1+\epsilon]$.
2.1. GRPO with Clipping
Note that in our context $x=\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}>0$ and the advantage $A(q,o)$ can be positive or negative and hence if $A(q,o)>0$ we have :
$f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)$ $=A(q,o)\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\text{clip}(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon,1+\epsilon)\right)$
$=A(q,o)\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)$
and if $A(q,o)<0$
$f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)$ $=A(q,o)\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},\text{clip}(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon,1+\epsilon)\right)$
$=A(q,o)\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)$
@@ 2.1. GRPO with Clipping
Recall that our reward $r$ is a verifiable reward that evaluates correctness of a reasoning or the execution of the code meaning that $r(q,o)\in\{0,1\}$. We note the probability of success of the old policy $\pi_{\text{old}}$:
$p:=p_{\theta_{\text{old}}}(q)=\mathbb{P}_{o\sim\pi_{\theta_{\text{old}}(.|q)}}(r(q,o)=1)$ (2)
Hence we have for mean and variance of a Bernoulli random variable :
$\mathbb{E}_{o^{\prime}\sim\pi_{\theta_{\text{old}}}(.|q)}r(q,o^{\prime})=p\text{ and }\mathsf{Var}_{o^{\prime}\sim\pi_{\theta_{\text{old}}}(.|q)}r(q,o^{\prime})=p(1-p).$
Assuming $10<p<$ 1 and replacing mean and variance in the advantage function (1) we obtain :
@@
$\begin{array}[]{cc}\displaystyle A(q,o)=\begin{cases}\frac{1-p}{\sqrt{p(1-p)}}&\text{if }r(q,o)=1,\\ -\frac{p}{\sqrt{p(1-p)}}&\text{if }r(q,o)=0.\end{cases}&\quad\text{equivalently}\quad\displaystyle A(q,o)=\begin{cases}\sqrt{\frac{1-p}{p}}&\text{if }r(q,o)=1,\\ -\sqrt{\frac{p}{1-p}}&\text{if }r(q,o)=0.\end{cases}\end{array}$
Hence we have conditioning on $q$ and using advantage expressions and $f_{\epsilon}$ for positive and negative advantage we obtain:
$\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)$ $=\sqrt{\frac{1-p}{p}}\,\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),\,r(q,o)=1}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)$
$-\sqrt{\frac{p}{1-p}}\,\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),\,r(q,o)=0}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)$
and hence the overall cost is obtained by taking expectation over $q$, note that $p=p_{\theta_{\text{old}}}(q)$:
@@
$-{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\sqrt{\frac{p_{\theta_{\text{old}}}(q)}{(1-p_{\theta_{\text{old}}}(q))}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)\mathbbm{1}_{r(q,o)=0}}-\beta\mathsf{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$
We see that GRPO is effectively a weighted contrastive loss that is weighted by ratio depending on the probability of succes of $\pi_{\theta_{\text{old}}}(.|q)$. We see from the weights plots that :
  • if the success probability of old policy is high ($p>0.5$), the weighting for points with success is low since the old policy is already good, and for failing point the weight is high and hence they are more penalized.
@@
  • if the success probability of old policy is low ($p<0.5$), the weighting for points with success is high since we want to reinforce those successes, and for failing points these are still penalized but with a small weight.
In summary, the standardized reward or the advantage function used in GRPO results in an interesting adaptive weighted contrastive loss : if the probability of success of the old policy is high, the wrong answers are more penalized than the correct ones are reinforced. If the probability of success of old policy is low, the correct answers are more reinforced than the wrong answers are penalized.
2.2. Stabilized GRPO with Clipping
@@
Note that in the previous section we assumed that $10<p<$ 1, we alleviate this in the following by adding a smoothing factor $\varepsilon\in(0,1]$ in the advantage as follows:
$A(q,o)=\frac{r(q,o)-\mu}{\sqrt{\sigma^{2}+\varepsilon}}.$
This results with the following advantage:
@@
$A(q,o)=\begin{cases}\frac{1-p}{\sqrt{p(1-p)}+\epsilon}&\text{if }r(q,o)=1,\\ -\frac{p}{\sqrt{p(1-p)}+\epsilon}&\text{if }r(q,o)=0.\end{cases}$ (3)
$\omega^{+}_{\varepsilon}(p)$ $=\frac{1-p}{\sqrt{p(1-p)+\varepsilon}}$
$\omega^{-}_{\varepsilon}(p)$ $=\frac{p}{\sqrt{p(1-p)+\varepsilon}}.$ (4)
@@
Replacing the stabilized advantage in Equation (GRPO-Clip), conditionally on a prompt $q$ we obtain the following contrastive loss:
$\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)\mathbbm{1}_{r(q,o)=1}+\ \mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A(q,o)\right)\mathbbm{1}_{r(q,o)=0}$
$=\omega^{+}_{\varepsilon}(p)\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)\mathbbm{1}_{r(q,o)=1}-\omega^{-}_{\varepsilon}(p)\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)\mathbbm{1}_{r(q,o)=0},$
which results in the following contrastive optimization :
@@
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\Big{\{}\omega^{+}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)\mathbbm{1}_{r(q,o)=1}...$
@@
.$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}....-\omega^{-}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)\mathbbm{1}_{r(q,o)=0}\Big{\}}-\beta\mathrm{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$
2.3. Stabilized GRPO with No Clipping
@@
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\Big{\{}\omega^{+}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}\mathbbm{1}_{r(q,o)=1}..$
..$~{}~{}..-\omega^{-}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}\mathbbm{1}_{r(q,o)=0}\Big{\}}-\beta\mathrm{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$ (GRPO-No-Clip)

### v4

@@ Abstract.
Group Relative Policy Optimization (GRPO) was introduced in (Shao et al., 2024) and used to train DeepSeek–R1 (Guo et al., 2025) for promoting reasoning in LLMs under verifiable (binary) rewards. We show that the mean+variance calibration of these rewards induces a contrastive loss in which the contrastive samples are synthetic data drawn from the previous policy. While GRPO was originally paired with clipping to keep updates near the old policy, we analyze variants that differ in reward normalization (mean-only vs. mean+variance) and in how they regularize updates using KL divergence: either penalizing divergence from the previous model (mirror), penalizing divergence from a fixed reference model $\pi_{\mathrm{ref}}$, or combining both forms of regularization. For each, the optimal policy $\pi_{n}$ admits an explicit form in terms of the binary reward and the first and second order statistics of the reward under $\pi_{n-1}$, as well as the policies $\pi_{n-1}$ and $\pi_{\mathrm{ref}}$. Iterating results in a sequence $\{\pi_{n}\}$ whose probability of success (PoS) obeys a simple recurrence that converges to a fixed point determined by the reference PoS and the regularization strength. We further show that this fixed point exceeds the reference, demonstrating that GRPO amplifies the policy’s probability of success.
@@ 1. Introduction
We focus in this paper on Reinforcement Learning with Verifiable Rewards (RLVR) using GRPO, as recently termed by Lambert et al. (2024). Verifiable rewards for RL with LLMs typically include (i) correctness checks via string matching to a gold answer when available or via an LLM-as-judge otherwise (Guo et al., 2025; Hugging Face, 2024; Luo et al., 2025; Guan et al., 2025). Additionally, (ii) execution-based pass/fail in code generation (interpreters and unit tests) and (iii) simple binary checks for formatting/refusals provide scalable 0/1 signals for training (Hugging Face, 2024; Guo et al., 2025; Lambert et al., 2024). Verifiable reward balance simplicity and bias and are thought to be less prone to reward hacking than reward models learned from preference data. We note that a recent paper (Vojnovic and Yun, 2025) studies GRPO with a focus on the policy obtained using an approximation of the $\mathsf{KL}$ divergence used in practical implementations.
The original GRPO’s practical recipe (Shao et al., 2024) combines PPO-style clipping with an explicit KL regularizer to a frozen reference model. On the other hand, mirror-descent style updates that regularize to the previous iterate (rather than a fixed reference) have been studied under the Mirror Descent Policy Optimization (MDPO) framework, which interprets each step as approximately solving a trust-region problem via a Bregman (KL) proximity term to $\pi_{n-1}$ (see for example (Schulman et al., 2015; Tomar et al., 2021; Gunter et al., 2024)). “Dr. GRPO” (Liu et al., 2025) is a variant that removes variance normalization (i.e., uses mean-only normalization of group rewards), simplifying the scaling while keeping the same overall training loop. Finally, recent large-scale systems such as DAPO (Yu et al., 2025) report strong results when removing the reference-model KL entirely (i.e., training reference-free), alongside additional engineering choices such as decoupled clipping and dynamic sampling.
@@
    Contrastive Loss (Sec. 2). We show that GRPO with calibrated verifiable rewards is equivalent to an adaptive, weighted contrastive loss evaluated on samples from the previous policy.
@@
    Policy Recursions. Leveraging this equivalence, we derive, for multiple GRPO variants, a closed-form recursion for the optimal policy as a function of $\pi_{\mathrm{ref}}$, $\pi_{n-1}$, and the previous policy’s probability of success (PoS) $p_{n-1}$. Section 3 analyzes GRPO (no clipping) with a KL penalty to the reference; Section 4 studies Mirror GRPO with a KL penalty to the previous iterate only; Appendix E covers the mixed (two-KL) case i.e mixed KL penalties to reference and previous iteration; and Section 5 treats the mean-only normalization.
@@
    PoS Dynamics & Fixed-Point Amplification. We prove that the induced PoS sequence $(p_{n})$ satisfies a recursion admitting a fixed point $p^{*}$ and, under mild assumptions, $p_{n}\to p^{*}$ with $p^{*}\geq p_{\mathrm{ref}}$, establishing success amplification for GRPO. The stepwise monotonicity of $(p_{n})$ depends on the specific variant. The dynamic of the PoS is verified empirically in Appendix A. Code is provided in supplementary material.
@@ 2. GRPO With verifiable Rewards as an Adaptive Weighted Contrastive Loss
We note the mean and variance of the reward under a policy $\nu$ as follows:
$\mu_{\nu}(q)=\mathbb{E}_{o^{\prime}\sim\nu(.|q)}r(q,o^{\prime})\quad\sigma^{2}_{\nu}(q)=\mathsf{Var}_{o^{\prime}\sim\nu(.|q)}r(q,o^{\prime}).$
@@
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A_{\pi_{\theta_{\text{old}}}}(q,o)\right)-\beta\mathsf{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$ (GRPO-Clip)
where the “advantage” for an outcome $o$, $A(q,o)$ is given by the whitened reward:
$A_{\pi_{\theta_{\text{old}}}}(q,o)=\frac{r(q,o)-\mu_{\pi_{\theta_{\text{old}}}}(q)}{\sigma_{\pi_{\theta_{\text{old}}}}(q)},$ (1)
and for $\epsilon\in[0,1]$, the clipping function $f_{\epsilon}$ is given by $f_{\epsilon}(x,y)=\min(xy,\text{clip}(x,1-\epsilon,1+\epsilon)y).$
@@
We see that GRPO optimizes the whitened reward (referred to as advantage, $A(q,o)$, in (Shao et al., 2024)) using importance sampling from the “old” policy while maintaining the optimized policy close to $\pi_{\mathrm{ref}}$ as measured by the $\mathsf{KL}$ divergence. The clipping used in (GRPO-Clip) ensures that the likelihood ratio between the policy and the old policy is maintained within $1-\epsilon,1+\epsilon]$.
2.1. Whitening the Rewards in GRPO As means of Calibration
@@
Recall that our reward $r$ is a verifiable reward that evaluates correctness of a reasoning or code execution, so $r(q,o)\in\{0,1\}$. We note the probability of success of the old policy $\pi_{\text{old}}$:
$p(q)=p_{\theta_{\text{old}}}(q)=\mathbb{P}_{o\sim\pi_{\theta_{\text{old}}(.|q)}}(r(q,o)=1)$ (2)
Hence, for a Bernoulli random variable, the mean and variance are::
$\mu_{\pi_{\theta_{\text{old}}}}(q)=p(q)\text{ and }\sigma^{2}_{\pi_{\theta_{\text{old}}}}(q)=p(q)(1-p(q)).$
Let us assume in the following that $0<p(q)<1$ so that $\sigma^{2}_{\pi_{\theta_{\text{old}}}}(q)>0$. Replacing mean and variance in the whitened reward in (1) we obtain :
@@
$\begin{array}[]{cc}\displaystyle A_{\pi_{\theta_{\text{old}}}}(q,o)=\begin{cases}\frac{1-p(q)}{\sqrt{p(q)(1-p(q))}}&\text{if }r(q,o)=1,\\ -\frac{p(q)}{\sqrt{p(q)(1-p(q))}}&\text{if }r(q,o)=0.\end{cases}&\text{i.e,}\quad\displaystyle A_{\pi_{\theta_{\text{old}}}}(q,o)=\begin{cases}\sqrt{\frac{1-p(q)}{p(q)}}&\text{if }r(q,o)=1,\\ -\sqrt{\frac{p(q)}{1-p(q)}}&\text{if }r(q,o)=0.\end{cases}\end{array}$ (3)
Calibrated reward behavior.
We see that the whitening or the normalization of the verifiable reward in GRPO, calibrates the reward with respect to the conditional distribution of the reward under $\pi_{\theta_{\text{old}}}(.|q)$ for every prompt $q$. This normalization results in a calibration of the reward that involves non linear functions of the probability of the success (PoS) of the old policy $p(q)$. See Figure 1 for an illustration. For a correct answer $r(q,o)=1$, the calibrated reward is positive and decreases with the PoS $p(q)$: rare successes (small $p(q)$) receive more credit than easy ones (large $p(q)$). For an incorrect answer ($r(q,o)=0$), the calibrated reward is negative, and its absolute value is increasing with $p(q)$. Wrong outcomes are more penalized when success is likely (for high $p(q)$) and less penalized when success is rare (low $p(q)$).
@@
2.2. GRPO with verifiable Reward As a Weighted Contrastive Loss
Replacing the calibrated reward $A_{\pi_{\theta_{\text{old}}}}(q,o)$ ((3)) for a verifiable reward in GRPO’s optimization objective (GRPO-Clip) we obtain:
$\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}f_{\epsilon}\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},A_{\pi_{\theta_{\text{old}}}}(q,o)\right)$ $=\sqrt{\frac{1-p(q)}{p(q)}}\,\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),\,r(q,o)=1}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)$
$-\sqrt{\frac{p(q)}{1-p(q)}}\,\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q),\,r(q,o)=0}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right),$
where we used that for $x>0$ and $y>0$, $f_{\epsilon}(x,y)=x\min(y,1+\epsilon)$ and for $x>0,y<0$, $f_{\epsilon}(x,y)=x\max(y,1-\epsilon)$.
The overall cost is further obtained by taking expectation over $q$, noting $p(q)=p_{\theta_{\text{old}}}(q)$:
@@ 2.2. GRPO with verifiable Reward As a Weighted Contrastive Loss
${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\sqrt{\frac{p_{\theta_{\text{old}}}(q)}{(1-p_{\theta_{\text{old}}}(q))}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)\mathbbm{1}_{r(q,o)=0}}-\beta\mathsf{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$
@@
We see that GRPO is effectively a weighted contrastive loss that is weighted by a ratio depending on the probability of success of $\pi_{\theta_{\text{old}}}(.|q)$. We see from the weights plots that if the success probability of the old policy is high ($p>0.5$), the weighting for points with success is low since the old policy is already good, and for failing points the weight is high and hence they are more penalized. On the other hand if the success probability of old policy is low ($p<0.5$), the weighting for points with success is high since we want to reinforce those successes, and for failing points these are still penalized but with a small weight.
2.3. Stabilized GRPO
@@
Note that in the previous sections we assumed that $0<p(q)<1$, so we ensure $\sigma^{2}_{\pi_{\theta_{\text{old}}}}(q)>0$. In the following, we alleviate this in the following by adding a smoothing factor $\varepsilon\in(0,1]$ in the advantage as follows:
$A_{\pi_{\theta_{\text{old}}}}(q,o)=\frac{r(q,o)-\mu_{\pi_{\theta_{\text{old}}}}(q)}{\sqrt{\sigma^{2}_{\pi_{\theta_{\text{old}}}}(q)+\varepsilon}}.$
This results with the following stabilized whitened reward:
@@
$A_{\pi_{\theta_{\text{old}}}}(q,o)=\begin{cases}+\omega^{+}_{\varepsilon}(p(q)),&r(q,o)=1,\\ -\omega^{-}_{\varepsilon}(p(q)),&r(q,o)=0,\end{cases}\quad\omega^{+}_{\varepsilon}(p)=\frac{1-p}{\sqrt{p(1-p)+\varepsilon}},\qquad\omega^{-}_{\varepsilon}(p)=\frac{p}{\sqrt{p(1-p)+\varepsilon}},$ (4)
with smoothing $\varepsilon>0$.
@@
Replacing the stabilized advantage in (GRPO-Clip), we obtain the following contrastive optimization:
@@
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\Big\{\omega^{+}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\min\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1+\epsilon\right)\mathbbm{1}_{r(q,o)=1}$
@@
$~~~~~~~~~~~~~~~~~~~~~~-\omega^{-}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\max\left(\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)},1-\epsilon\right)\mathbbm{1}_{r(q,o)=0}\Big\}-\beta\mathrm{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$
Stabilized GRPO with No Clipping
To simplify Equation GRPO-Clip, let us consider this objective without the clipping ($\epsilon\to+\infty$); we obtain:
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}A_{\pi_{\theta_{\text{old}}}}(q,o)-\beta\mathsf{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$ (GRPO)
@@
$\max_{\theta}\mathbb{E}_{q\sim\rho_{\mathcal{Q}}}\Big\{\omega^{+}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}\mathbbm{1}_{r(q,o)=1}$
$-\omega^{-}_{\varepsilon}(p_{\theta_{\text{old}}}(q))\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}(.|q)}\frac{\pi_{\theta}(o|q)}{\pi_{\theta_{\text{old}}}(o|q)}\mathbbm{1}_{r(q,o)=0}\Big\}-\beta\mathrm{KL}(\pi_{\theta}||\pi_{\mathrm{ref}})$ (GRPO-No-Clip)